OpenStudy (anonymous):

Find the polynomial f(x) that has the roots of -2, 3 of multiplicity 2. Explain how you would verify the zeros of f(x).

OpenStudy (campbell_st):

so it means that the the function is f(x) = (x + 2)(x -3)^2 if a root is x = -2 then the factor is (x + 2) x = 3 then the factor is (x - 3) multiplicity of 2 means the factor is squared. so the general form is \[f(x) = a(x +2)(x -3)^2\] to test it, select a point on the curve..... substitute x and y and then you'll find the left is a multiple of the right.... then multiple is the value of a

OpenStudy (anonymous):

Thanks you!!! :D

Directrix (directrix):

Does the prepositional phrase "of multiplicity 2" describe the root 3 only or does the phrase describe BOTH roots 3 and -2?

Directrix (directrix):

If both, then we would have this--> f(x) = (x + 2)^2 (x -3)^2

OpenStudy (anonymous):

Okay thanks!!! :D)

Directrix (directrix):

Enjoyed working on the Descartes problem.